### v1

# A self-supervised framework for learning whole slide representations

Abstract
Whole slide imaging is fundamental to biomedical microscopy and computational pathology. However, whole slide images (WSIs) present a complex computer vision challenge due to their gigapixel size, diverse histopathologic features, spatial heterogeneity, and limited/absent data annotations. These challenges highlight that supervised training alone can result in suboptimal whole slide representations. Self-supervised representation learning can achieve high-quality WSI visual feature learning for downstream diagnostic tasks, such as cancer diagnosis or molecular genetic prediction. Here, we present a general self-supervised whole slide learning (S3L) framework for gigapixel-scale self-supervision of WSIs. S3L combines data transformation strategies from transformer-based vision and language modeling into a single unified framework to generate paired views for self-supervision. S3L leverages the inherent regional heterogeneity, histologic feature variability, and information redundancy within WSIs to learn high-quality whole-slide representations. We benchmark S3L visual representations on two diagnostic tasks for two biomedical microscopy modalities. S3L significantly outperforms WSI baselines for cancer diagnosis and genetic mutation prediction. Additionally, S3L achieves good performance using both in-domain and out-of-distribution patch encoders, demonstrating good flexibility and generalizability.
Self-supervised learning, representation learning, histopathology, biomedical microscopy
Figure 1: Self-supervised whole slide learning. Previous work in computational pathology relies on multiple instance learning with weak supervision from slide or patient level labels to learn whole slide representations (Zhu et al., 2017; Ilse et al., 2018; Lu et al., 2021b; Shao et al., 2021; Chen et al., 2022b; Javed et al., 2022; Chen et al., 2023). Here, we present a self-supervised framework for learning whole slide representations, called S3L, by combining data transformations from vision and language modeling to generate high-quality paired views.
1 Introduction
Whole slide imaging is an integral part of pathology and laboratory medicine. Computational pathology can provide rapid tissue analysis, such as tumor detection, subtyping, and grading, of complex WSIs. Modern computational methods have advanced beyond morphology-based diagnostics to “omics” predictions, prognostication, and treatment response predictions from WSIs alone (Hollon et al., 2023; Coudray et al., 2018; Chen et al., 2022c). Previous results have relied on pathologic annotations for whole slide representation learning; unfortunately, these annotations are weak and expensive to obtain (Campanella et al., 2019), resulting in suboptimal whole slide representations. Few previous studies have focused on learning high-quality whole slide representations using self-supervision (Chen et al., 2022b, 2023; Lazard et al., 2023). Self-supervised whole slide learning can enable better generalization performance to a wide range of downstream pathology tasks with minimal to no annotations required.
In this paper, we present a general and scalable self-supervised framework for whole slide representation learning, S3L. S3L treats gigapixel WSIs as a sequence of patch tokens and applies a domain-informed set of vision-language transformations, including splitting, cropping, and masking, to generate high-quality views for self-supervised training. Over a range of state-of-the-art patch encoders, S3L learns high-quality patch feature aggregation and whole slide representations compared to feature pooling and non-transformer baselines. The main contributions are:
  • We introduce S3L, a general, flexible, and lightweight SSL framework for WSIs at scale and benchmark on two WSI domains and tasks.
  • We design and evaluate the domain-informed vision-language transformations for WSI view generation.
  • We demonstrate that S3L achieves good performance using both in-domain and out-of-distribution patch encoders.
Figure 2: S3L overview. A. The S3L framework consists of a two-stage model architecture: 1) a pre-trained patch encoder $\mathcal{E}$ and 2) a transformer whole slide encoder $f$. WSIs are first divided into small patches, and the patch encoder extracts patch-level features. We then apply whole slide transformations to the patch tokens to create two views of the same WSI. The transformations combine splitting, cropping, and masking, and are informed by the structure and unique properties of WSIs. Each augmented views are encoded by the transformer whole slide encoder, and the feature learning can be self-supervised by any SSL paradigm. B. Example SSL paradigms. In our experiments, we focus on three representative SSL paradigms, including SimCLR, BYOL, and VICReg.
2 Related Work
2.1 Computational pathology
Computational pathology combines whole slide imaging and computer vision methods for the analysis of biomedical tissues. Cancer diagnosis, prognostication, and response-to-treatment are some of the most common computer vision tasks within computational pathology. WSIs pose a unique computer vision challenge due to image sizes ranging up to 150K $\times$ 150K pixels. Data annotations are often limited to whole slide or patient-level labels (Campanella et al., 2019). Moreover, WSIs have a different inherent structure compared to natural images, such as being non-object-centric, containing regional heterogeneity, and visual feature redundancy.
Despite these challenges, significant progress has been made over the last decade due to more accessible slide scanners and modern deep learning methods. Cancer diagnosis (Campanella et al., 2019; Lu et al., 2021a, b; Hollon et al., 2020), molecular classification (Coudray et al., 2018; Hollon et al., 2023), and prognostication (Chen et al., 2021) with WSIs represent how computational pathology can be used to directly improve patient care. Whole slide representation learning for gigapixel WSI search and public WSI-natural language supervision represent future directions in computational pathology (Chen et al., 2022a; Huang et al., 2023).
2.2 Multiple instance learning
Multiple instance learning (MIL) is a type of supervised learning task where the labels are available only for bags of instances, rather than each individual instance (Dietterich et al., 1997). Learning to classify WSIs has generally been regarded as a MIL task, where each patch is an instance and the WSI is a bag of patches (Ilse et al., 2018). There are two main components in MIL for WSI: the patch encoder and aggregator. Early works directly train the patch encoder in a weakly supervised fashion with WSI labels and generate a whole slide representation with mean pooling (Zhu et al., 2017; Hou et al., 2016). Many works (Campanella et al., 2019; Li et al., 2019; Lu et al., 2021b; Shao et al., 2021) used a ResNet model pretrained on ImageNet as their patch encoder, while there were alternative patch encoder architectures such as MI-FCN (Yao et al., 2019, 2020). (Ilse et al., 2018) first used an attention-based aggregator for WSI classification, and most later works used variations of attention or more powerful transformer-based aggregation models (Li et al., 2019; Yao et al., 2020; Li et al., 2021a; Lu et al., 2021b; Shao et al., 2021; Wang et al., 2022b; Yu et al., 2023; Lin et al., 2023). MCAT (Chen et al., 2021) combined WSI with genomic information for multimodal classification. Additive MIL (Javed et al., 2022) introduced intrinsic interpretability to an attention-based aggregator.
2.3 Self-supervised representation learning
Self-supervised learning (SSL) is a method of learning representations that does not require labels or annotations. SSL defines pretext learning tasks, such as instance discrimination or reconstruction, as training objectives and is typically evaluated by the performance of the learned representations on downstream tasks (Jaiswal et al., 2020). SSL has shown success in natural images (Chen et al., 2020; Grill et al., 2020; Bardes et al., 2021), natural language processing (Devlin et al., 2018; Mikolov et al., 2013), and multimodal learning (Lu et al., 2023; Huang et al., 2023). SSL has been used for patch encoding in WSIs. DSMIL (Li et al., 2021a) applied SimCLR to learn patch representations. HIPT (Chen et al., 2023) built a general-purpose patch encoder for all types of H&E patches and tasks with self-supervised DINOv2 training. Most recently, PLIP (Huang et al., 2023) leverages histology images and comments posted to Twitter for self-supervised vision-language pretraining.
While there have been attempts at learning whole slide representations in an end-to-end manner with supervision (Hemati et al., 2021), or enhancing whole slide performance with a supervised contrastive loss (Wang et al., 2022a), few works have focused on learning self-supervision beyond patches. HiDisc (Jiang et al., 2023) uses a data hierarchy to learn patch representations that are invariant to slide and patient membership. HIPT (Chen et al., 2022b) uses a self-supervised patch encoder and an encoder for relatively small regions, but weak supervision is used for learning WSI representations.
Giga-SSL (Lazard et al., 2023) and SS-MIL (Tavolara et al., 2022) attempted to learn whole slide representations using self-supervision by masking patch tokens. Both methods have shown limited success compared to their baselines. Giga-SSL also emulates slide augmentation using pre-computed augmented patch representations. However, this strategy does not span all combinations of possible random augmentations, is memory/compute intensive, and is infeasible for large-scale datasets.
3 Methods
3.1 The S3L framework
WSIs are partitioned into smaller non-overlapping fields-of-view, or patches. Let $\boldsymbol{I}=\left\{I_{i}\right\}_{i=1}^{n}$ be a WSI, where $n$ is the number of patches in a WSI, and each $I_{i}\in\mathbb{R}^{3\times H\times W}$, with coordinate $p_{i}\in\mathbb{Z}^{2}$. $\boldsymbol{p}=\left\{p_{i}\right\}_{i=1}^{n}$ are the coordinates of all patches in the WSI. S3L is a two-stage learning framework to learn WSI representations: 1) a pre-trained patch encoder $\mathcal{E}$; and 2) a transformer-based whole-slide encoder $f$. The pre-trained patch encoder $\mathcal{E}$ can have the architecture of any visual feature extractor. The overall model architecture of S3L is illustrated in Figure 2.
In the S3L framework, the patch encoder $\mathcal{E}$ encodes each $I_{i}$ into a patch token, ${x}_{i}=\mathcal{E}(I_{i})$. We represent all patch tokens in a WSI as $\boldsymbol{x}=\{x_{i}\}_{i=1}^{n}=\mathcal{E}(\boldsymbol{I})$. The whole-slide encoder $f$ serves as an aggregation function that learns a whole-slide representation using patch tokens and their corresponding coordinates: $h=f(\boldsymbol{x},\boldsymbol{p})=f(\mathcal{E}(\boldsymbol{I}),\boldsymbol{p})$. $f$ can be any learned aggregation architecture, such as in ABMIL (Ilse et al., 2018) or a transformer (Vaswani et al., 2017; Shao et al., 2021). Due to the scale of the WSI, it is infeasible to both train the patch encoder $\mathcal{E}$ and whole-slide encoder $f$ jointly in an end-to-end manner. Thus, we freeze the patch encoder $\mathcal{E}$ to allow for large mini-batch training of the whole-slide encoder $f$. The aim of the S3L framework is to learn high-quality whole slide representations with $f$ using self-supervision.
Siamese self-supervised methods share a common strategy: apply two random transformations to a single data example to generate two distinct views, called a positive pair. We designed our S3L framework to be compatible with existing siamese SSL objectives, where a WSI undergoes a transformation into different views: $t(\boldsymbol{I})=[I_{i}\dots I_{k}]$, $t(\boldsymbol{p})=[p_{i}\dots p_{k}]$, $t^{\prime}(\boldsymbol{I})=[I_{j}\dots I_{\ell}]$, and $t^{\prime}(\boldsymbol{p})=[p_{j}\dots p_{\ell}]$, where $t,t^{\prime}\in\mathcal{T}$ are randomly drawn from a set of transformations described in section 3.2. The patches and their corresponding coordinates from these transformed views are processed through $\mathcal{E}$ and $f$ to obtain the whole slide representations:
$h=f\left(\mathcal{E}(t(\boldsymbol{I})),t(\boldsymbol{p})\right),\quad h^{\prime}=f\left(\mathcal{E}(t^{\prime}(\boldsymbol{I})),t^{\prime}(\boldsymbol{p})\right).$
An SSL loss is used to self-supervise whole slide training:
$\mathcal{L}_{\text{S3L}}\left(h,h^{\prime}\right).$
In our experiments, we selected representative methods from different SSL families as identified by (Balestriero et al., 2023), including SimCLR (Chen et al., 2020) from contrastive learning, BYOL (Grill et al., 2020) from self-distillation, and VICReg (Bardes et al., 2021) from canonical component analysis.
3.2 Whole slide transformations
While previous work focused on pixel space transformations (Lazard et al., 2023), we hypothesized that patch augmentations are alone insufficient to generate high-quality views for whole slide SSL. A patch encoder trained with instance discrimination, for example, should be invariant to pixel-space augmentations. Thus, pixel space augmentations, such as random cropping, should have minimal effect on the representation space given a strong patch encoder. Similar representations across augmentations are especially significant when the patch encoder is trained with a self-supervised loss, as illustrated with SimCLR in Figure 3, because it is explicitly enforced as a pretext task. We opt to bypass patch augmentations altogether, thereby reducing S3L memory and compute burden, and focus on transformations at the WSI level.
Figure 3: Limited effect of pixel-level patch augmentations. We qualitatively evaluate the effect of pixel-level augmentation on the patch representations by visualizing the tSNE plot of SimCLR pretrained patch representations sampled from a single WSI. We observe that strong augmentations at pixel-level have a minimal effect on the patch embeddings. The invariant behavior of the patch encoder is explicitly enforced by the SimCLR pretext task.
Figure 4: S3L transformation strategy. S3L transformation strategy combines splitting, cropping, and masking, and they are informed by the size, region diversity, and information redundancy of WSIs. Splitting partitions patches into mutually exclusive sets decreases mutual information between views; cropping can generate spatially diverse views covering different regions on the WSI; masking reduces redundant visual features and improves training efficiency. The combination of these transformations can create optimal positive pairs for whole slide representation learning.
WSI are gigapixel images divided into a sequence of tokens. Therefore, our whole slide transformations combine strategies from both vision and language modeling, which are selected to address the domain-specific properties of WSI (Figure 4): splitting, cropping, and masking.
  • Splitting to decrease mutual information between views. Reducing mutual information between SSL views, while keeping task-relevant information intact, improves downstream performance (Tian et al., 2020). The splitting transformation randomly partitions patch tokens into disjoint sets, ensuring mutually exclusive set membership for each token. Enforcing mutual exclusion reduces mutual information between SSL views.
  • Cropping to capture regional heterogeneity. WSIs contain regional differences in histologic features. Similar to multi-cropping in visual SSL (Chen et al., 2020; Caron et al., 2020, 2021), cropping generates spatially diverse views of WSI with variable sizes and features.
  • Masking to reduce redundant visual features. Due to their gigapixel size, WSIs can contain large regions of similar histologic features and tissue phenotypes, causing redundancy. Masking is an effective transformation for vision and language modeling (Devlin et al., 2018; He et al., 2022; Chen et al., 2022b). For each view, the masking transformation samples a subset of patch tokens without replacement. Additionally, masking increases training efficiency by reducing sequence lengths.
While splitting, cropping and masking can be used individually, we apply them sequentially to generate versatile and high-quality views for S3L training.
3.3 S3L Implementation
End-to-end whole slide learning with large batch sizes is infeasible due to the gigapixel scale. To enable efficient training, all unaugmented patch tokens were stored prior to whole slide training, and S3L transformations were applied to patch tokens. Each whole slide was represented as a embedding $\boldsymbol{x}\in\mathbb{R}^{n\times d}$ and a coordinate $\boldsymbol{p}\in\mathbb{Z}^{n\times 2}$. Splitting and masking were implemented as row-wise partitioning and dropout on $\boldsymbol{x},\ \boldsymbol{p}$, respectively. Cropping was implemented as coordinate filtering on $\boldsymbol{p}$. A detailed description of our implementation with pseudocode is in Appendix B.
4 Experiments
We evaluated our method on two previously benchmarked clinical pathology modalities and tasks: intraoperative brain tumor classification using stimulated Raman histology (SRH) and molecular classification of diffuse gliomas using hematoxylin and eosin (H&E) staining.
S3L was compared to pooling baselines, and to a supervised method, SupCon (Khosla et al., 2020), as the performance upper bound. In addition, we evaluated S3L qualitatively through visualization of attention heatmaps. We evaluated S3L’s robustness and generalizability by training with different patch encoders.
4.1 Datasets and Tasks
Stimulated Raman histology classification
We evaluated S3L on a multiclass histopathological brain tumor diagnosis task using an SRH dataset. SRH is an optical microscopy method that provides rapid imaging of unprocessed biological tissues (Freudiger et al., 2008; Orringer et al., 2017), optimally suited for intraoperative applications. Our SRH dataset was collected at the University of Michigan, and the study has been approved by the Institutional Review Board (HUM00083059). Our imaging protocol follows (Jiang et al., 2022), with informed consent from each patient. The SRH dataset consists of an annotated and an unlabeled set. The annotated set includes 2960 WSIs labeled by board-certified expert pathologists and consists of normal brain tissue and six common brain tumor diagnoses: high-grade glioma, low-grade glioma, meningioma, pituitary adenoma, schwannoma, and metastatic tumor. This annotated set was designed for supervised training and evaluation, and was divided into training and evaluation sets with 2035 and 925 slides, respectively. The unlabeled set encompassing 2919 WSIs was intended for self-supervised pretraining in conjunction with the annotated training set. Detailed slide breakdown for each class can be found in Appendix A.
H&E molecular classification
We also evaluated S3L using publicly available brain tumor H&E stained WSIs from The Cancer Genome Atlas (TCGA) and Digital Brain Tumour Atlas (DBTA) (Roetzer-Pejrimovsky et al., 2022). We focused on the molecular classification of diffuse gliomas, the most common malignant brain tumor. Our dataset consisted of three diffuse glioma subgroups as defined by the World Health Organization (Louis et al., 2021): glioblastoma, astrocytoma, and oligodendroglioma. Molecular classification is a challenging computer vision task because the diagnoses are made via molecular testing (such as genetic sequencing), and are not possible for expert pathologists using H&E images. Our H&E dataset comprises a total of 2650 WSIs from both datasets, with 2309 training slides and 341 slides from the TCGA dataset set for evaluation. H&E Slide composition can be found in Appendix A.
4.2 Implementation details
We used a ResNet-34 (He et al., 2016) patch encoder to extract 512-dimension features for all patches in each slide. The patch encoder was pre-trained with SimCLR (Chen et al., 2020) and HiDisc (Jiang et al., 2023) for both datasets and was frozen during whole slide training. We used a six-layer transformer (Vaswani et al., 2017) slide encoder with the same dimension as patch tokens and 4 heads, with learnable Fourier features positional embeddings (Li et al., 2021b) for better spatial representation. Learned positional embeddings are visualized in Appendix B.2. We followed common practice for most SSL methods and use a two-layer projection head. For BYOL, we used a one-layer projection head and a one-layer prediction head. We used an AdamW optimizer and a cosine decay scheduler after warm-up in the initial 10% of the iterations. The learn rate was adjusted between $1\times 10^{-3}$ and $1\times 10^{-7}$ to accommodate the unique training dynamics of different SSL methods and tasks. With our slide transformation strategy, we adjusted cropping and masking sizes for each experiment and utilize up to 64 patches per slide for each view. SRH and H&E experiments were conducted over 300 and 400 epochs, respectively, with an effective batch size of up to 1024 WSIs. All experiments were done using one NVIDIA A40 GPU with mixed-precision training. Training with the SimCLR and VICReg losses takes approximately 6 hours and 8 hours for BYOL. SupCon experiments are trained for 100 epochs.
4.3 Evaluation Protocol
Since linear and finetuning evaluation protocols are sensitive to hyperparameters, such as learning rate, we used a $k$ nearest neighbor ($k$NN) classifier for self-supervised evaluation (Caron et al., 2021; He et al., 2020). For each model, we extracted the whole slide representation for both training and testing data by using all the patches in each slide. The nearest neighbor classifier was then used to match each whole slide representation in the testing dataset to the $k$ most similar whole slide representation in the training dataset based on their cosine similarity. H&E evaluation used only WSI from TCGA training set. We report mean class accuracy (MCA) and F1 scores since both datasets are imbalanced, and standard deviation is computed across three random seeds for each experiment.
5 Results
We first evaluated the S3L frameworks with different families of SSL objectives. We then compared the generalizability of S3L on different state-of-the-art and out-of-domain patch encoders. We present the ablation studies for S3L and discuss the role of each transformation. Finally, we visualize the self-attention coefficients on the CLS token from the last layer of the whole slide encoder.
5.1 S3L benchmark with different SSL objectives
We benchmark S3L with patch encoder pretrained with SimCLR. Max and mean feature pooling are non-parametric aggregation methods with strong performance for good patch encoders (Ilse et al., 2018; Chen et al., 2022b; Li et al., 2021a; Shao et al., 2021). As shown in Table 1, S3L surpasses max and mean pooling across all self-supervised objectives. Best S3L performance approaches supervised SCM performance upper bound, with an MCA difference of only 2.7% for SRH and 4.0% for H&E. We included an alternative supervised method, CLAM (Lu et al., 2021b), a non-transformer-based SOTA model for whole slide classification. Whole slide SCM transformations with SupCon outperform CLAM by +1.5%. These results demonstrate the flexibility and performance boost of SCM transformations for both SSL and supervised training.
SRH H&E
MCA F1 MCA F1
Max pooling 73.5 75.6 68.2 69.2
Mean pooling 72.5 73.2 68.3 67.9
S3L-SimCLR 77.3 (0.6) 77.6 (0.3) 76.5 (0.5) 76.1 (0.4)
S3L-VICReg 82.3 (0.5) 82.3 (0.3) 72.8 (0.7) 72.4 (0.9)
S3L-BYOL 79.4 (0.4) 78.7 (1.2) 74.6 (0.9) 73.8 (0.9)
CLAM 83.5 (1.3) 82.9 (1.2) 78.9 (2.2) 77.7 (1.8)
Supervised SCM 85.0 (0.5) 84.2 (0.6) 80.5 (1.6) 79.9 (1.4)
Table 1: S3L performance on SRH and H&E self-supervised benchmarks. SimCLR patch tokens are used for training and evaluation. For pooling, mean/max pooling operation is performed over all patch tokens for each slide, and uses kNN evaluation. S3L with SimCLR loss is represented as S3L-SimCLR, for example. Each experiment is run 3 times and a mean (std) is reported.
5.2 Out-of-Distribution generalization of S3L
We explored the generalization ability of S3L with a variety of patch encoders with different model architectures, training algorithms, and dataset domains. A list of SOTA patch encoders is described in Table 2. Based on their training data, patch encoders were grouped into three categories: Out-of-Distribution (OOD), near-domain, and in-domain. ImageNet patch encoder was considered OOD for both SRH and H&E because of the domain gap between WSI and natural images. HIPT (Chen et al., 2022b) was trained with H&E WSIs from TCGA. PLIP (Huang et al., 2023) was trained with images and comments from Twitter. Both include data from multiple organ systems and institutions. Therefore, they were considered near-domain for our brain tumor H&E benchmark and OOD for our SRH benchmark. The SimCLR patch encoder, as in Section 5.1, and HiDisc patch encoder are in-domain since they use the training data from our benchmarks. The performance of S3L with the SimCLR loss is shown in Figure 5, with extended metrics available in Appendix C.
Patch
encoder
Architecture
Training
dataset
Training
algorithm
ImageNet ResNet34 ImageNet Cross entropy
HIPT ViT-S/16 All of TCGA DINO
PLIP ViT-B/32 OpenPath CLIP
SimCLR ResNet34 In-domain (see 4.1) SimCLR
HiDisc ResNet34 In-domain (see 4.1) HiDisc
Table 2: Patch encoders. We benchmark S3L with different patch encoders. ImageNet, HIPT, and PLIP patch encoders use off-the-shelf models trained on OOD or near-domain datasets. SimCLR and HiDisc patch encoders are trained with in-domain data.
Figure 5: S3L performance with different patch encoders. For SRH, ImageNet, HIPT, and PLIP patch encoders are pre-trained on OOD data. For H&E, HIPT and PLIP patch encoders are considered near-domain.
S3L compliments patch encoders.
For the SRH benchmark, the patch encoders trained on OOD data fail to generalize, as demonstrated by the poor pooling performance shown in Figure 5. However, the patch encoders learn general visual features that were leveraged by S3L. After training with S3L, WSI representations improved dramatically, as demonstrated by a maximum 18.2% boost in MCA for ImageNet. For the H&E benchmark, S3L compensated for OOD and near-domain patch encoders, surpassing in-domain pooling performance. In addition, S3L closes the performance gap between OOD, near-domain, and in-domain patch encoders, all being within 4% MCA. Overall, these results demonstrate that S3L can enhance whole-slide representation learning using encoders from different domains, thereby reducing the reliance on a specialized in-domain patch encoder. S3L training time ranges from 6-8 GPU hours, while the patch encoder is computationally intensive and it usually takes over 80 GPU hours to train. These results highlight S3L’s versatility and flexibility with different patch encoders and its broader impacts on efficient training for WSI representations.
S3L boosts performance of in-domain patch encoders.
Among the patch encoders, in-domain HiDisc is the highest performing, as demonstrated by the MCA metric for both benchmarks, regardless of the aggregation method. HiDisc patch encoder enforces slide discrimination as an inductive bias during patch training as part of its learning objective. S3L boosts whole slide representations even in the setting of high-quality, in-domain patch encoders. We found that S3L shows more improvement in H&E than the SRH benchmark. This could be due to higher variability between different regions within each H&E WSI. S3L allows the transformer to encode a slide representation from multiple patch tokens, instead of using slide membership of single patches to learn patch representations, as employed in HiDisc.
5.3 Ablation studies
These ablation experiments are conducted under the same configurations unless otherwise noted.
Figure 6: Attention heatmap visualization. Self-attention visualization demonstrates S3L’s ability to distinguish between different tissue phenotypes such as blood, dense tumor, and necrosis. To our knowledge, this is the first self-supervised transformer-based attention maps on full gigapixel WSIs, instead of localized regions.
Transformation choices
Table 3 reports the performance of individual and combinations of slide-level transformations. We leave out splitting alone since it does not reduce the total number of patch tokens for both views, making it prohibitive to train. For the masking and cropping alone, we show that cropping performs better than masking for both SRH and H&E. Combining cropping and masking shows similar performance with cropping alone, but is more efficient to train due to the reduced number of tokens per view. Combining each transformation with splitting improves or maintains performance, showing the benefit of further decreasing mutual information between views. Overall, cropping is the most important transformation, because capturing regional heterogeneity is the most effective and challenging pre-text task.
SRH H&E
MCA F1 MCA F1
M 75.4 (0.9) 76.0 (0.7) 69.9 (0.5) 69.7 (0.5)
C 81.3 (1.5) 81.5 (1.5) 74.3 (0.2) 74.4 (0.2)
S+M 76.1 (0.3) 76.3 (0.8) 69.2 (1.5) 69.3 (1.5)
S+C 82.1 (1.4) 82.2 (1.2) 73.6 (0.6) 73.8 (0.4)
C+M 81.2 (0.5) 81.8 (0.5) 75.5 (2.1) 75.4 (2.2)
S+C+M 82.3 (0.5) 82.3 (0.3) 76.5 (0.5) 76.1 (0.4)
Table 3: Ablation over transformations. S, splitting, C, cropping, M, masking.
Transformation parameters
We further examine the relationship between mutual information and model performance by combining and adjusting the slide transformations. We studied the effect of maximizing the number of tokens included in each view, with SCM and masking only. The results are illustrated in Figure 7. Mutual information levels can be adjusted using SCM, and these transformations should be used in conjunction with each other to optimize model performance. Previous studies have concluded that fewer patch tokens resulted in better model performance (Lazard et al., 2023), but only in the case of using masking transformation alone. Additional transformation ablation studies with crop area sizes are in Appendix C.3.
Figure 7: Max token limit ablation. SCM, splitting-cropping-masking. Shaded area, std across three different random seeds.
5.4 Self-attention visualizations reveal tissue phenotypes in full gigapixel WSIs
We evaluate S3L qualitatively by visualizing the self-attention on the full WSI. As illustrated in Figure 6, self-attention maps on H&E WSI can distinguish different tissue phenotypes such as blood, dense tumor, and necrosis. The CLS token attends to dense tumor regions instead of non-diagnostic regions such as blood and necrosis for slide representations. This observation is consistent with existing literature on self-supervised vision transformers, where attention maps can serve as unsupervised segmentation (Caron et al., 2021; Chen et al., 2022b). To our knowledge, our work is the first to generate self-supervised transformer-based attention maps on full gigapixel WSIs, instead of localized regions, showing the strong representation capacity, flexibility, and efficiency of S3L. More H&E and SRH attention map visualizations are in Appendix C.4.
6 Conclusion
We present S3L, a transformer-based SSL approach for WSI representation learning, in a general, versatile, and lightweight framework. S3L uses domain-informed, vision-language transformations for high-quality WSI view generation and self-supervised learning. We demonstrated that S3L achieved superior performance on two computational pathology benchmarks and datasets, cancer diagnosis and molecular genetic prediction. To our knowledge, this is the first systematic study of whole slide representation learning evaluating the role of SSL objectives, transformation strategies, and patch encoders.
Limitations
Our experimentation is limited to two biomedical microscopy modalities and three major self-supervised learning paradigms. Future work includes validating S3L on additional datasets, including H&E from different organ systems, or fluorescent microscopy, as well as different SSL paradigms such as generative SSL. Another limitation of S3L is hyperparameter tuning for different configurations of SCM transformation, which are domain and dataset-specific. The future direction of our work also includes efficient end-to-end training to unify the training of patch and slide encoders.
Impact Statement
This paper presents with the aim to advance the field of whole slide representation learning and computational pathology. Our work builds toward an automated diagnostic support system for biomedical microscopy and histopathology, as it paves the way for more accurate diagnoses and personalized treatment recommendations. Moreover, we strive to develop self-supervised learning frameworks in order to promote and maintain data privacy and ethical data use of patient data within healthcare AI systems. Additionally, SSL is essential for developing medical foundation models. We hope that S3L can serve as a potential foundation model training strategy within computational pathology. While our experiments are in medical applications, it has a potential impact on other similar gigapixel image modalities, such as geospatial and astronomical images.
Acknowledgement
We would like to thank Karen Eddy and Lin Wang for their administrative support and data collection.
This work is supported in part by grants NIH T32GM141746, NIH F31NS135973, NIH K12NS080223, NIH R01CA226527, Cook Family Brain Tumor Research Fund, Mark Trauner Brain Research Fund: Zenkel Family Foundation, Ian’s Friends Foundation, and the Investigators Awards grant program of Precision Health at the University of Michigan.
T.H. is a shareholder of Invenio Imaging, Inc., a company developing SRH microscopy systems.
Appendix A Data Details
SRH dataset.
Our SRH benchmark includes common brain tumors and normal brain tissue histology. Our data collection and processing follows the protocol described in (Jiang et al., 2022), where each specimen is scanned with a commercially available stimulated Raman histology microscope manufactured by Invenio Imaging. Each specimen is diagnosed by a board-certified neuropathologist and a detailed number of slides per class for SRH is in Table 4.
Class Training Validation
HGG 407 132
LGG 210 107
Mening. 434 204
Metast. 236 114
Pit. 448 194
Schwan. 47 22
Normal 253 152
Unlabeled 2919 -
Table 4: SRH dataset breakdown. Number of slides in the training and validation set for supervised training and KNN evaluation. HGG, high grade glioma, LGG, low grade glioma, mening., meningioma, metast., metastasis, pit., pituitary adenoma, schwan., schwannoma, normal, normal brain tissue.
H&E dataset.
Our H&E benchmark includes glioma specimens from the Cancer Genome Atlas Program (TCGA) and the Digital Brain Tumour Atlas (DBTA) (Roetzer-Pejrimovsky et al., 2022). These publicly available datasets feature H&E images collected in both the United States and Europe. Slides are divided into three classes based on molecular labels as defined by the World Health Organization (Louis et al., 2021), using IDH and 1p/19q co-deletion status included in both datasets. At training time, both DBTA and TCGA training set is used, and a separate, held-out validation set from the TCGA dataset is used for benchmarking. A detailed number of slides per class is in Table 5. Our H&E data processing pipeline follows (Jiang et al., 2023). Each WSI is divided into 300x300 patches, and blank / background regions are excluded. All patches used in training are stain-normalized using the Macenko algorithm (Macenko et al., 2009).
Class Molecular label Training Validation
DBTA TCGA TCGA
Oligo. IDH-mut, 1p/19q-codel 176 265 62
Astro. IDH-mut 157 360 88
GBM IDH-wt 619 732 191
Table 5: H&E dataset breakdown. Number of slides in the training and validation set from both TCGA and DBTA. Oligo, oligodendroglioma, Astro, astrocytoma, GBM, glioblastoma.
Appendix B Methods and Implementation
B.1 SCM transformation details
As described in section 3.3, we implement our SCM transformations with pre-computed embeddings and patch coordinates. Recall that each whole slide is represented as a $(\boldsymbol{x},\boldsymbol{p})$ tuple, where $\boldsymbol{x}\in\mathbb{R}^{n\times d}$ is the embedding tensor and $\boldsymbol{p}\in\mathbb{Z}^{n\times 2}$ is a corresponding coordinate tensor for each patch in the WSI. Each transformation is implemented by indexing into the rows (corresponding to patches/tokens) of $\boldsymbol{x}$ and $\boldsymbol{p}$. A PyTorch style pseudocode is in Algorithm 1.
Splitting.
Splitting has one hyperparameter: the ratio of tokens between two views. The ratio specifies the fraction of tokens in the first view and is used to compute the number of tokens in each view. Tokens are split randomly between views accordingly, and two disjoint views are returned.
Cropping.
Cropping has two sets of hyperparameters: cropping area range, and cropping aspect ratio range. The cropping area and cropping aspect ratio are randomly chosen from their respective ranges. A random token in the view with coordinates $(r_{A},c_{A})$ is selected to be an anchor as the crop center of the view. A coordinate range $(r_{0},c_{0}),(r_{1},c_{1})$ is computed based on the area and aspect ratio:
$H$ $=\sqrt{\text{Area}/\text{aspect}}$
$W$ $=H\cdot\text{aspect}$
$r_{0}$ $=r_{A}-H/2;$ $r_{1}$ $=r_{A}+H/2$
$c_{0}$ $=c_{A}-W/2;$ $c_{1}$ $=c_{A}+W/2,$
where $H,W$ are the height and width of the crop. Tokens in this coordinate range are included in the transformed view.
Masking.
[…]
Algorithm 1 S3L transformations in PyTorch style
# hyperparameters:
# - split_ratio: fraction of token in first view
# - (area_min, area_max): cropping area range
# - (aspect_min, aspect_max): cropping aspect ratio range
# - (mask_min, mask_max): masking ratio range
# - max_token_lim: max masking token limit
@torch.no_grad()
def transform(x, p):
    # inputs:
    # - x: patch embeddings (n d)
    # - p: patch coordinates (n 2)
    # output:
    # - (x1, p1): patch embs and coords for the first view
    # - (x2, p2): patch embs and coords for the second view
    (x1, p1), (x2, p2) = split(x, p)
    (x1, p1) = mask(crop(x1, p1))
    (x2, p2) = mask(crop(x2, p2))
    return (x1, p1), (x2, p2)
def split(x, p):
    # randomly permute index of all tokens
    rand_idx = torch.randperm(len(p))
    # determine index for each view
    view1_sz = int(split_ratio * len(p))
    idx1, idx2 = rand_idx[:view1_sz], rand_idx[view1_sz:]
    # filter tokens for each view
    return ((x[idx1,:], p[idx1,:]), (x[idx2,:], p[idx2,:]))
def crop(x, p):
    # randomly choose an anchor as center of the crop
    anchor_id = torch.randint(high=len(p))
    anchor = p[anchor_id, :]
    # randomly select crop area and aspect ratio
    crop_area = torch.randint(low=crop_min, high=crop_max)
    aspect = torch.FloatTensor(1).uniform_(aspect_min, aspect_max)
    # compute min and max coordinate of the crop
    height = torch.sqrt(areas / aspect)
    width = (height * aspect)
    r0, r1 = anchor[0] - height / 2, anchor[0] + height / 2
    c0, c1 = anchor[1] - width / 2, anchor[1] + width / 2
    # filter tokens
    idx = ((p[:,0] > r0) & (p[:,0] < r1) &
           (p[:,1] > c0) & (p[:,1] < c1))
    return x[idx,:], p[idx,:]
def mask(x, p):
    # randomly select a number of tokens to keep
    mask_ratio = torch.randint(low=mask_min, high=mask_max)
    size = mask_ratio * len(p)
    # optional: set a upper bound on number of tokens
    size = torch.minimum(size, max_token_lim)
    # randomly generate index and filter tokens
    idx = torch.randperm(len(p))[:sizes]
    return x[idx,:], p[idx,:]
B.2 Relative positional embedding
To handle the different sizes of WSIs, it is non-trivial to use a fixed-size learnable position embedding with absolute distance (Devlin et al., 2018). We adopt the relative positional encoding introduced in (Li et al., 2021b), which utilizes the learnable Fourier feature, modulated with a multi-layer perceptron. This positional encoding takes the coordinates of the patch encoder as input and the output is added directly to the transformer with the same dimension. We visualize the similarity of positional embeddings in Figure 8.
Figure 8: Cosine similarity of learned positional embeddings. Visualization is generated similarly as in (Dosovitskiy et al., 2020): each patch shows the cosine similarity between the position embedding of the token with the indicated row and column and the position embeddings of all other tokens.
Appendix C Extended Results
In this section, we present extended results presented in section 5.
C.1 S3L results with ImageNet and HiDisc patch encoders
In Figure 5, we compared the S3L’s performance to the mean pooling baseline and supervised SCM performance upper bound. We present the detailed benchmark results for ImageNet and HiDisc patch encoders in Tables 6 and 7, respectively. S3L significantly improves slide representations in all self-supervised learning paradigms experimented, for both ImageNet and HiDisc patch encoders. Notably, in our H&E benchmark, S3L slide representations trained with the ImageNet patch encoder outperform pooling results with the HiDisc patch encoder, demonstrating S3L’s versatility in learning slide representations with out-of-distribution patch encoders.
SRH H&E
MCA F1 MCA F1
Max pooling 53.5 53.7 57.4 57.6
Mean pooling 53.1 53.0 59.8 74.2
S3L-SimCLR 71.7 (0.7) 73.2 (0.4) 74.0 (1.9) 73.0 (1.9)
S3L-VICReg 66.2 (0.7) 66.6 (0.7) 71.7 (0.7) 71.0 (1.1)
S3L-BYOL 69.3 (1.6) 69.9 (1.2) 75.3 (2.5) 74.0 (2.5)
Supervised SCM 75.4 (1.4) 76.1 (0.8) 76.4 (1.1) 76.4 (1.2)
Table 6: S3L results with ImageNet patch encoder. Each experiment is run 3 times, and a mean (standard deviation) is reported. SCM, splitting-cropping-masking, MCA, mean class accuracy, F1, F1 score.
SRH H&E
MCA F1 MCA F1
Max pooling 82.3 83.5 73.3 73.4
Mean pooling 82.7 82.9 74.9 60.1
S3L-SimCLR 83.4 (0.1) 82.9 (0.3) 77.5 (1.3) 77.1 (1.3)
S3L-VICReg 84.0 (0.7) 83.9 (0.7) 75.3 (1.7) 75.0 (1.7)
S3L-BYOL 84.2 (1.4) 83.9 (1.2) 75.4 (2.1) 75.0 (2.0)
Supervised SCM 87.4 (0.8) 87.5 (0.7) 83.6 (0.3) 82.4 (0.2)
Table 7: S3L results with HiDisc patch encoder. Each experiment is run 3 times and a mean (standard deviation) is reported. SCM, splitting-cropping-masking, MCA, mean class accuracy, F1, F1 score.
Patch encoder Slide encoder SRH H&E
MCA F1 MCA F1
HIPT Max pooling 47.5 47.2 58.5 59.0
Mean pooling 55.1 56.4 54.7 55.0
S3L-SimCLR 70.1 (0.4) 71.9 (0.5) 74.9 (1.7) 73.9 (1.7)
Supervised SCM 72.5 (1.1) 73.7 (0.9) 76.6 (1.8) 76.1 (1.6)
PLIP Max pooling 50.2 50.4 63.3 62.6
Mean pooling 51.7 51.9 66.8 66.3
S3L-SimCLR 68.2 (0.4) 69.7 (0.4) 76.2 (2.0) 75.5 (2.2)
Supervised SCM 72.1 (1.4) 72.6 (1.2) 82.6 (0.7) 81.5 (0.9)
Table 8: Out of distribution S3L benchmarks. Each experiment is run 3 times and a mean (standard deviation) is reported. SCM, splitting-cropping-masking, MCA, mean class accuracy, F1, F1 score.
C.2 S3L results with out-of-distribution patch encoders
In Figure 5, we also benchmarked the S3L’s performance when trained with HIPT (Chen et al., 2022b) and PLIP (Huang et al., 2023) patch encoders. We present the detailed benchmarks for these experiments with F1 score in Table 8. Recall that the HIPT patch encoder is trained with the complete TCGA dataset, and the PLIP patch encoder is trained with publicly available (non-WSI) histology images on Twitter. For SRH, HIPT and PLIP patch encoders are OOD, and for H&E, HIPT and PLIP are near domain.
C.3 Extended ablation results
We present additional ablation experiments varying cropping size range in Figure 9. Cropping size is a hyperparameter that allows us to adjust the amount of mutual information between views, i.e. larger crop size $\rightarrow$ higher probability and area of overlapping regions between views. Our ablation study shows that for SimCLR training on our H&E benchmark, $[100,400]$ is the optimal cropping size range, achieving a MCA of 76.5. These results correspond to the hypothesis in (Tian et al., 2020), where an optimal amount of mutual information between views achieves the best model performance.
Figure 9: Cropping size range ablation for H&E. Cropping size range is on the x-axis, where numbers in the bracket correspond to the min and max cropping size. None, no cropping applied, Shaded area, standard deviation across three different random seeds.
C.4 Extended attention heatmaps
We present extended attention heatmap visualization for SRH and H&E in Figures 10 and 11, respectively. Recall these attentions are the self-attention of the CLS token from the last layer of the whole slide encoder. On both SRH and H&E heatmaps, we can see that the whole slide encoder attends to clinically significant regions such as dense tumors, and regions with tumor infiltration, and avoids non-diagnostic regions such as blood, tissue processing artifacts, and blood vessels.
Figure 10: Extended attention heatmap for H&E whole slides. The attention map shows the S3L-trained whole slide transformer can differentiate between different morphologies like dense tumors, blood, and different artifacts on the gigapixel WSIs. Visualization shows high attention to varying degrees of tumor infiltration, and low attention to low cellularity, blood, necrotic regions, and tissue processing artifacts. Oligo, oligodendroglioma, Astro, astrocytoma, GBM, glioblastoma.
Figure 11: Extended attention heatmap for SRH whole slides. On this SRH WSI, relatively small compared to H&E, the model shows a strong capacity for unsupervised segmentation of histological features. Visualization shows high attention values to tumor regions, and low attention values to non-diagnostic regions such as blood vessels, laser noise, and empty space.